In each of the following cases, find the unknown variable: (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Account Break Even Unit Price Unit Variable Vost Fixed Cost Depreciation
125,736 $39 $30 $820,000 ?
165,000 ? 27 2,320,000 975,000
21,430 92 ? 237,000 128,700
Answer #1

Account Break Even=(Fixed Cost+Depreciation)/(Unit price-Unit Variable cost)

Account Break Even Unit price Unit Variable cost Fixed Cost Depreciation
125,736 39 30 820,000 [125,736*(39-30)-820,000]=$311,624
165,000 [27+(2,320,000+975,000)/165,000]=46.97(Approx). 27 2,320,000 975,000
21430 92 [92-(237,000+128,700)/21430]=74.94 237,000 128700